OpenStudy (anonymous):

Mayank, Mirza, Little and Jagbir bought a motorbike for $60. Mayank contributed half of the total amount contributed by others. Mirza contributed one – third of total amount contributed by others, and little contributed one – fourth of the total amount contributed by others. What was the money paid by Jagbir?

OpenStudy (anonymous):

Mayank = $20 Mirza = $15 Little = $12 Jagbir = $13

OpenStudy (anonymous):

set up a system of equations let Mayank be x, Mirza be y, little be z, Jagbir be j, and total be t: x=1/2(y+z+j) y=1/3(x+z+j) z=1/4(x+y+j) t=x+y+z+j set the first three equal to zero and the last one equal to t which is 60. Solve the system using a TI-89 or similar is the easy way. I got: x=20 y=15 z=12 j=13(Jagbir)

OpenStudy (anonymous):

Half of the rest equals a third. A third of the rest equals a quarter. A quarter of the rest equals a fifth.